# Dataforge: A Data Agent Platform for Autonomous Data Engineering

### v1

@@
Abstract
The growing demand for AI applications in fields such as materials discovery, molecular modeling, and climate science has made data preparation an important but labor-intensive step. Raw data from diverse sources must be cleaned, normalized, and transformed to become AI-ready, while effective feature transformation and selection are essential for efficient training and inference. To address the challenges of scalability and expertise dependence, we present Data Agent, a fully autonomous system specialized for tabular data. Leveraging large language model (LLM) reasoning and grounded validation, Data Agent automatically performs data cleaning, hierarchical routing, and feature-level optimization through dual feedback loops. It embodies three core principles: automatic, safe, and non-expert friendly, which ensure end-to-end reliability without human supervision. This demo showcases the first practical realization of an autonomous Data Agent, illustrating how raw data can be transformed “From Data to Better Data.”
I Introduction
@@
Modern AI applications increasingly rely on AI-ready data—data that are clean, consistent, and semantically aligned with model objectives. In fields such as materials discovery, molecular design, and climate modeling, AI-ready data determine whether models can achieve reliable prediction, efficient training, and meaningful scientific insight [yu2025ai]. Yet, transforming raw data into such AI-ready form remains a persistent challenge. Raw datasets are often heterogeneous, incomplete, and noisy, collected from instruments or repositories with inconsistent formats and semantics. Converting them into usable inputs requires extensive cleaning, normalization, and feature transformation, which are still largely manual, error-prone, and dependent on domain expertise. This growing mismatch between the rapid progress of AI models and the slow, fragmented process of data preparation exposes a fundamental data engineering gap that increasingly limits the scalability and reliability of AI applications.
@@
The rapid progress of large language models (LLMs) has opened a new path toward autonomous data engineering. LLMs not only capture extensive domain knowledge from large-scale text corpora but also exhibit emerging abilities to interpret data schemas, infer transformation logic, and generate executable operations—tasks traditionally reserved for human experts. Although many recent works have explored using LLMs for data-centric tasks such as LLM-based feature engineering and transformation [han2024large, hollmann2023large, horn2019autofeat, zhang2023openfe, gong2025evolutionary], they remain limited in several aspects: (1) Lack of full automation: Most approaches still require user intervention, manual code review, or step-by-step prompting to complete a data pipeline. (2) Limited reliability: LLMs may generate invalid or unsafe operations without schema verification or grounding checks, leading to broken workflows. (3) Expert dependency: Many systems assume the user understands LLM prompts, data schemas, or tool APIs, which prevents non-expert adoption.
@@
Newer LLMs now demonstrate emergent agentic behaviors—the ability to reason, plan, and act through iterative interaction with tools and environments. This shift marks an agentic evolution in how intelligence is deployed: systems no longer just generate outputs for a descriptive query, but autonomously understand tasks, reason, plan, ground, execute actions, and adapt through feedback. Extending this paradigm to the data domain gives rise to the concept of Data Agents: autonomous systems that can understand datasets, reason about transformations, and execute data-centric operations to improve quality, consistency, and utility.
@@
Figure 1: Conceptual comparison between a traditional manual workflow and the agentic workflow.
Figure 2: The framework of the Dataforge system.
@@
To this end, we develop the Dataforge, an autonomous data agent specialized for tabular data that embodies three core design goals: 1) Automatic: performs end-to-end data transformation without manual intervention; 2) Safe: ensures workflow reliability via routing, grounding, and execution validation; 3) Non-expert Friendly: allows users with minimal technical background to improve their data by simply specifying task and data types. The dataagent is implemented based on the perception–planning–grounding–execution principles [fu2025autonomous] as a deployable system for tabular data. The data agent includes a sequential stages of data cleaning, hierarchical routing, grounding validation, and dual-loop optimization. In particular, the hierarchical routing mechanism determines both the high-level task family and the low-level feature actions; each action is checked before execution to ensure safety, and results are evaluated to guide iterative improvement; the dual feedback loops ensure that each decision is validated before execution and that performance improvements are optimized through experience replay. Our Contribution: present the first functional data agent demo that leverages LLMs to automatically reason over tabular data and iteratively orchestrate next grounded actions with autostop to fabric AI-ready data.
II Dataforge: Agentic Data Engineering
II-A Overview of the Dataforge
Fig. 2 shows that Dataforge aims to automate the end-to-end data preparation process—from raw input to adaptive refinement—through a structured six-stage pipeline. (1) Input and Cleaning: Raw data is standardized via schema alignment, missing-value handling, and type correction under preset modes (light, aggressive, or time-series). (2) Task-Level Routing: A rule-based router identifies the task type (e.g., classification, clustering, regression) and dispatches data to corresponding action-level LLM router. (3) Action-Level Routing: Within each branch, an LLM-based planner selects, transforms, or generates features based on metadata and experience. (4) Grounding and Execution: Actions are validated for schema and logic safety, then executed through modular Python or SQL engines. (5) Evaluation and Feedback: An evaluator computes task-specific metrics (e.g., F1, AUC, RMSE) and maintains two feedback loops to regulate the process, in which the inner loop validates actions to prevent workflow failure, and the outer loop optimizes performance across iterations. (6) Experience Replay and Summary: The system logs all operations as experiences to replay for continual learning and automatically generate reports for traceable provenance.
II-B Hierarchical Routing
After data cleaning, to enable efficient and reliable decision-making, we adopt a hierarchical routing architecture, including task-level and action-level reasoning. At the task-level routing, a rule-based router quickly identifies the task type: classification, regression, or unsupervised learning, based on table schema metadata, such as, data types, label structures, and feature distribution. Such lightweight router relies on deterministic heuristics, instead of large language models, thus, enable fast and reliable responses across diverse datasets. At the action-level routing, a compact LLM-based planner refines the decision by selects and plans the most suitable feature-level actions such as, different ordered combinations of feature selection, transformation, or generation, under the identified task (e.g., a classification dataset). Since each router operates within a smaller, well-defined action space, this hierarchical routing approach not only accelerates processing but also avoid invalid or high-risk operations.
II-C Dual Feedback Loops
We develop two collaborative feedback loops to transform the static workflow into an adaptive, self-correcting process, in order to achieve autonomy and continual refinement. 1) Action Validation Loop for Safety: This feddback loop is to ground actions to ensure operational safety before execution. Each planned action is first grounded through schema alignment, type checking, and logical consistency tests, such as, detecting divisions by zero or invalid type conversions. Only actions that pass validation proceed to execution so as to prevent runtime errors and maintaining workflow integrity. 2) Performance Optimization Loop for Adaptiveness: This feedback loop focuses on post-execution improvement and enabling the Dataforge to converge toward optimal behavior without human supervision. In particular, after each valid action runs, an evaluator measures task-specific metrics (e.g., F1, AUC, RMSE). If performance improves, the system updates its optimal configuration; otherwise, it triggers re-planning of new data engineering actions. The two loops establish a closed-cycle control mechanism for both grounded safety and self-optimizing adaptivity.
II-D Automatic Summary and Reporting
@@
To ensure transparency and provenance, the system automatically log the routing choices, all actions, grounding outcomes, machine learning performance metrics and summarize these logs into a structured concise natural language report that outlines: (i) the cleaning and feature operations applied, (ii) the evolution of routing decisions, (iii) the validation checks performed, and (iv) the performance progression across iterations. The summary, viewable through the demo interface, translates internal reasoning into a human-readable narrative that lets non-experts trace data transformations and decisions, but also allows the feedback loop to collect experiences as benchmark data and perform self-reflection.
II-E User Interface
Fig. 3 illustrates the user interface of Dataforge, which offers an intuitive environment for autonomous tabular data transformation. Users can upload datasets in standard formats (.csv, .parquet, and .hdf), label predictors and response column in a machine learning task, and configure runtime parameters, such as, Max Actions, Min Improvement, and Time Budget. After uploading datasets, users can select from preloaded examples. During execution, the interface visualizes real-time logs of thinking steps in reasoning, data cleaning and feature engineering actions, and predictive performance metrics (e.g., F1-score, AUC, MAE, and RMSE depending on classification, regression, and unsupervised tasks ), so that users can observe how the data agent progressively refines data quality. Upon completion, it summarizes all executed actions, validation outcomes, and final performance statistics, to provide transparency and provenance, as well as collect action historical trajectories for benchmark data of training data agents. The interface is designed to be modular and extensible, supporting future integration with interactive dashboards and collaborative agent environments.
Figure 3: The interface of Dataforge.
III Demonstration Scenarios
III-A Demonstration Setup
All demonstrations are conducted on Ubuntu 22.04.3 LTS with an Intel(R) Core(TM) i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB, CUDA 12.2). The system is implemented in Python 3.12 using PyTorch for GPU-accelerated data transformation.
III-B Quantitative Evaluations
Study of Overall Performances. TABLE I summarizes the overall performance of Dataforge compared with representative baselines, including classical AutoML methods, RL-based pipelines, and pure LLM approaches. Across all nine datasets, Dataforge achieves the highest or comparable predictive performance while maintaining full automation and zero training cost. Its rule-based task routing and LLM-driven action selection allow it to adapt to different task types efficiently, avoiding unstable or invalid operations common in purely generative systems.
Notably, the agent demonstrates consistent improvements over the baseline data across multiple datasets, validating its ability to autonomously refine data representations through feature transformation and selection. The performance gains are achieved with minimal user input—only specifying the data file, task type, and optional runtime constraints—highlighting the practicality of the proposed autonomous data-processing paradigm.

### v2

# Dataforge: Agentic Platform for Autonomous Data Engineering

@@
Abstract.
The growing demand for artificial intelligence (AI) applications in materials discovery, molecular modeling, and climate science has made data preparation a critical but labor-intensive bottleneck. Raw data from diverse sources must be cleaned, normalized, and transformed to become AI-ready, where effective feature transformation and selection are essential for robust learning. We present Dataforge, an LLM-powered agentic data engineering platform for tabular data that is automatic, safe, and non-expert friendly. It autonomously performs data cleaning and iteratively optimizes feature operations under a budgeted feedback loop with automatic stopping. Across tabular benchmarks, it achieves the best overall downstream performance; ablations further confirm the roles of routing/iterative refinement and grounding in accuracy and reliability. Dataforge demonstrates a practical path toward autonomous data agents that transform raw data from data to better data.
Data Agent, LLMs, Data-Centric Agentic AI
copyright: none
1. Introduction
@@
Modern AI applications increasingly rely on AI-ready data that are clean, consistent, and semantically aligned with the objectives of the model. In fields such as materials discovery, molecular design, and climate modeling, AI-ready data determine whether models can achieve reliable prediction, efficient training, and meaningful scientific insight (Yu, 2025). However, transforming raw data into such an AI-ready form remains a persistent challenge (Wang et al., 2025b; Ying et al., 2025c). Raw datasets are often heterogeneous, incomplete, and noisy, collected from instruments or repositories with inconsistent formats and semantics. Converting them into usable inputs requires extensive cleaning, normalization, and feature transformation, which are still largely manual, error-prone, and dependent on domain expertise. This growing mismatch between the rapid progress of AI models and the slow and fragmented data preparation process exposes a fundamental data engineering gap that increasingly limits the scalability and reliability of AI applications.
@@
The rapid progress of large language models (LLMs) has opened a new path toward autonomous data engineering. LLMs not only capture extensive domain knowledge from large-scale text corpora but also exhibit emerging abilities to interpret data schemas, infer transformation logic, and generate executable operations—tasks traditionally reserved for human experts (Xie and Ju, 2025). Although many recent works have explored using LLMs for data-centric tasks such as LLM-based feature engineering and transformation (Han et al., 2024; Hollmann et al., 2023; Horn et al., 2019; Zhang et al., 2023; Gong et al., 2025c), they remain limited in several aspects: (1) Lack of full automation: Most approaches still require user intervention, manual code review, or step-by-step prompting to complete a data pipeline. (2) Limited reliability: LLMs may generate invalid or unsafe operations without schema verification or grounding checks, leading to broken workflows. (3) Expert dependency: Many systems assume the user understands LLM prompts, data schemas, or tool APIs, which prevents non-expert adoption.
@@
Newer LLMs now demonstrate emergent agentic behaviors: the ability to reason, plan, and act through iterative interaction with tools and environments. This shift marks an agentic evolution in how intelligence is deployed: systems no longer just generate outputs for a descriptive query, but autonomously understand tasks, reason, plan, ground, execute actions, and adapt through feedback. Extending this paradigm to the data domain gives rise to the concept of Data Agents: autonomous systems that can understand datasets, reason about transformations, and execute data-centric operations to improve quality, consistency, and utility.
@@
Figure 1. Conceptual comparison between a traditional manual workflow and the agentic workflow.
@@
To this end, we develop the Dataforge, an autonomous data agent platform specialized for tabular data that embodies three core design goals: 1) Automatic: performs end-to-end data transformation without manual intervention; 2) Safe: ensures workflow reliability via routing, grounding, and execution validation; 3) Non-expert Friendly: allows users with minimal technical background to improve their data by simply specifying task and data types. The Dataforge is implemented based on the planning–grounding–execution principles (Fu et al., 2025) as a deployable system for tabular data. The data agent includes a sequential set of stages of data cleaning, hierarchical routing, grounding validation, and dual-loop optimization. In particular, the hierarchical routing mechanism determines both the high-level task family and the low-level feature actions; each action is checked before execution to ensure safety, and results are evaluated to guide iterative improvement; the dual feedback loops ensure that each decision is validated before execution and that performance improvements are optimized through experience replay.
Our empirical study demonstrates that Dataforge provides strong and robust gains across diverse tabular datasets: it achieves the best overall downstream utility on six classification and three regression benchmarks compared with classical, RL-based, and LLM-based approaches. Ablation results further verify that hierarchical routing and evaluation-driven refinement are key to performance, while grounding substantially improves executability by reducing invalid actions and repeated retries. It also avoids RL-style training cost and achieves strong performance with low call complexity, supporting its practicality as an automated data engineering tool.
Our Contribution: We present Dataforge, an LLM-powered agentic data engineering platform for tabular data that is automatic, safe, and non-expert friendly, iteratively orchestrating feature operations with automatic stopping to produce AI-ready data.
2. Preliminaries
2.1. Problem Formulation
We consider automated data preparation and feature engineering for tabular datasets. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a dataset with a feature table $X\in\mathbb{R}^{n\times d}$ (mixed numeric/categorical/time columns) and an optional target column $y$. Given $\mathcal{D}$ and a task specification $\mathcal{T}$ (e.g., classification or regression, evaluation metric, and optional constraints), our goal is to construct a transformation pipeline $\pi$ that maps the raw dataset to an AI-ready representation:
$\pi:\mathcal{D}\rightarrow\mathcal{D}^{\prime}.$
A pipeline $\pi$ is a sequence of executable feature operations $\pi=[a_{1},\ldots,a_{K}]$, where each action $a_{k}$ is selected from a tool library $\mathcal{A}$ (e.g., missing-value imputation, encoding, scaling, feature generation/selection). Applying $\pi$ produces a transformed dataset $\mathcal{D}^{\prime}$ and an execution log. We optimize $\pi$ under a user-defined budget $\mathcal{B}$ (e.g., time limit, maximum number of actions) to improve downstream performance while preserving correctness and safety:
$\max_{\pi}\;\mathrm{Perf}(\mathcal{D}^{\prime},\mathcal{T})\quad\text{s.t.}\quad\mathrm{Safe}(\pi,\mathcal{D})=1,\;\;\mathrm{Cost}(\pi)\leq\mathcal{B}.$
Here, $\mathrm{Perf}(\cdot)$ is measured by a standard metric (e.g., F1 for classification, 1-RAE for regression) using a fixed evaluation protocol, and $\mathrm{Safe}(\cdot)$ indicates whether the pipeline satisfies grounding checks.
2.2. Platform Value
End-to-end automation: Given a dataset and a minimal task specification, Dataforge automatically profiles data, proposes feature operations, validates them, executes transformations, and iteratively improves data utility under a budget. Reliability and safety: Unlike prompt-only approaches, Dataforge enforces action grounding and guardrails to prevent invalid or unsafe transformations and to ensure stable execution. Benchmarkable and reproducible outputs. Dataforge exports a replayable pipeline together with structured logs and summary reports. Non-expert usability: Dataforge exposes task-level controls rather than requiring users to craft prompts or write feature code, lowering the barrier for non-expert practitioners.
2.3. System I/O Contract
Inputs: (i) a tabular dataset (CSV/Parquet or a database table), (ii) an optional target column name, (iii) a task specification $\mathcal{T}$ including task type and evaluation metric, and (iv) optional constraints $\mathcal{B}$ such as time budget, maximum number of actions, and minimum improvement threshold for early stopping.
Outputs: (i) a transformed dataset $\mathcal{D}^{\prime}$ (AI-ready table), (ii) a replayable pipeline $\pi$ that specifies the ordered list of feature operations with parameters, (iii) execution logs that record dataset diffs and validation outcomes for each step, and (iv) a summary report that explains key transformations and the measured utility improvements.
Failure handling: If an action cannot be validated or executed, Dataforge returns a failure reason (e.g., schema mismatch, type violation, suspected leakage, or runtime error) and either (a) replans with an alternative action, or (b) terminates with the best valid pipeline found so far under the budget.
2.4. Important Concepts
AI-ready data: We use AI-ready to describe tabular data that are schema-consistent, type-valid, and suit to model training (e.g., reduced missingness and improved downstream performance).
Operation Set: To refine the feature space, we need to apply mathematical operations to existing features to generate new informative features. All operations are collected in an operation set, denoted by $\mathcal{O}$. These operations can be classified as unary and binary operations. The unary operations such as “square”, “exp”, “log”, etc. The binary operations such as “plus”, “multiply”, “minus”, etc.
Feature Transformation Sequence: Assuming a dataset $\mathcal{D}=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$ includes the original feature set $\mathcal{F}_{0}=\{f_{1},\ldots,f_{K}\}$ and predictive targets $y$. We transform the existing features using mathematical compositions $\tau$ consisting of feature ID tokens and operations to generate new and informative features. $K$ compositions are adopted to refine $\mathcal{F}_{0}$ to a better feature space $\tilde{\mathcal{F}_{0}}=\{\tilde{f}_{1},\ldots,\tilde{f_{K}}\}$. The collection of the $K$ compositions refers to the feature transformation sequence, which is denoted by $\Gamma=[\tau_{1},\cdots,\tau_{K}]$.
Postfix Expressions: The transformation sequence should be in a computable and machine-learnable format. The original infix representation has issues like redundancy, semantic sparsity, a high likelihood of illegal transformations, and an overly large search space. We introduce postfix expressions to solve these problems. Postfix expressions don’t need many brackets to determine calculation priority. Scanning from left to right suffices to reconstruct the corresponding sequence, greatly reducing sequence-modeling difficulty and computational cost. They also reduce the ambiguity of the transformation sequence. Most importantly, it reduces the search space from exponential to a finite set $|C|=|\mathcal{O}|+|\mathcal{F}_{0}|D+3$. Here, $|\mathcal{O}|$ represents the operation set size, $|\mathcal{F}_{0}|$ is the original feature set dimension, $D$ is feature numbers, and $3$ refers to start tokens <SOS><SOS>, separation token <SEP><SEP>, and end token <EOS><EOS>.
Figure 2. The framework of the Dataforge system.
3. Dataforge: Agentic Data Engineering
3.1. Overview of the Dataforge
Figure. 2 overviews Dataforge, which automates tabular data preparation and feature engineering via a structured, iterative pipeline. (1) Input and Cleaning: the input table is standardized through basic cleaning routines (schema/type normalization and missing-value handling) under preset modes (light, aggressive, or time-series). (2) Task-Level Routing: a lightweight router infers the learning setting (e.g., classification, regression, or unsupervised) and activates the corresponding feature-action space. (3) Action-Level Planning: conditioned on dataset profiles and past trajectories, an LLM planner proposes candidate feature operations (selection, transformation, and generation) from a tool library. (4) Grounding and Execution: each proposed action is validated by grounding checks (e.g., schema/type compatibility and logical constraints) before being executed by the backend. (5) Evaluation and Feedback: the evaluator measures task-specific metric (e.g., F1/AUC/RMSE) and drives two feedback loops: an inner loop rejects unsafe or invalid actions, while an outer loop performs iterative re-planning toward improved performance under a budget. (6) Experience and Reporting: the system records action trajectories, validation outcomes, and metric progression, enabling experience-guided planning and generating a human-readable report for traceability.
3.2. Hierarchical Routing
After data cleaning, Dataforge adopts a hierarchical routing strategy (Wang et al., 2025d; Bai et al., 2025a) to make decisions more accurate and efficient. Instead of directly searching over a large space of feature operations, we first perform a coarse-grained routing to a feature-operation family, and then conduct fine-grained planning within the selected family.
Task-level (coarse) routing: Here “task” refers to the data-engineering sub-task at the current iteration, i.e., choosing among three high-level operation families: feature selection, feature engineering, and feature transformation (Figure. 2). The router is rule-based and uses lightweight dataset profiles (e.g., feature dimensionality, type composition, missingness, cardinality, and simple distributional statistics) to decide which family is most promising to apply next. Intuitively, selection is preferred when the dataset exhibits strong redundancy or high dimensionality; transformation is preferred when scale/shape issues (e.g., heavy tails, outliers, or heterogeneous feature scales) dominate; and engineering is preferred when informative interactions or structured patterns are likely to exist. This coarse routing restricts subsequent planning to a smaller and more relevant candidate set, reducing confusion across operation types.
Action-level (fine) planning: Conditioned on the chosen family, an LLM-based planner selects concrete actions (e.g., specific selection criteria, transformation functions, or feature-generation methods) from the corresponding sub-library. By operating within a well-defined action subspace, the planner produces more consistent candidates and avoids irrelevant proposals, which improves routing accuracy and reduces unnecessary trials.
Why hierarchical routing: Let $\mathcal{A}$ denote the full tool library and $\mathcal{A}^{sel}$, $\mathcal{A}^{eng}$, $\mathcal{A}^{trans}$ denote the family-specific subsets. Hierarchical routing first chooses one subset and then searches within it, effectively shrinking the candidate space at each step. In our iterative setting, the outer optimization loop can revise the coarse choice in later iterations, enabling the system to alternate between families as needed.
3.3. Feature Engineering Tool Library
Dataforge organizes its feature operations into a family-structured tool library: feature selection, feature transformation, and feature engineering. Each tool corresponds to a parameterized feature action that can be proposed by the planner.
Feature selection: This family targets redundancy reduction and robustness by selecting informative subsets of columns. It includes standard filter and embedded operators (e.g., L1/LASSO, and tree-based importance), as well as group-wise selection primitives when features exhibit correlated blocks. Beyond classical operators, the library includes policy-based selectors inspired by reinforcement-learning-based selection (Liu et al., 2021a; Fan et al., 2021; Zhao et al., 2020; Wang et al., 2025e), which propose candidate subsets under budgeted evaluations.
Feature transformation: This family transforms existing features to improve numerical conditioning, handle scale and distribution mismatch, and mitigate noise/outliers, while preserving the original information content. We include common transformations such as normalization/standardization, power/log transforms, clipping for outliers, and discretization/binning. In addition to these hand-designed operators, the library can incorporate learned transformation modules that produce compact transformation programs or representations, leveraging reinforcement- or generation-based approaches to search over continuous or structured transformation spaces (Wang et al., 2023; Ying et al., 2024c; Gong et al., 2025d; Wang et al., 2022).
Feature engineering: This family constructs new features to expand representational capacity beyond the raw columns. We support widely used engineering primitives, including interaction/cross features (e.g., pairwise products and conjunctions), logical compositions, and group-wise aggregation patterns, which are effective when predictive signals emerge from multi-column interactions. For structured domains, the library can further incorporate domain-aware generators such as categorical hashing representations with hierarchical crossing and topology-aware feature space reconstruction for graph-structured attributes (Ying et al., 2025a, 2023), enabling systematic feature creation when hand-crafted templates are insufficient.
3.4. Dual Feedback Loops
We develop two collaborative feedback loops that turn the workflow into an adaptive, self-correcting process. The first loop ensures that every proposed action is executable (grounding for safety), while the second loop guides the system toward better downstream performance through iterative evaluation and re-planning.
Action Validation Loop for Safety: Given a candidate action proposed by the planner, Dataforge performs grounding checks before execution. These checks validate (i) schema compatibility (target columns exist and the action’s output schema is well-formed), (ii) type consistency (e.g., numeric operators are not applied to categorical strings; encoders receive valid category sets), and (iii) logical feasibility (e.g., avoiding division by zero, invalid logarithms, or degenerate transformations that collapse a column to constants). If any check fails, the action is rejected with an explicit failure reason, and the planner is prompted to revise the proposal. Only grounded actions proceed to execution, which significantly reduces runtime errors and prevents invalid pipelines from contaminating subsequent iterations.
Performance Optimization Loop for Adaptiveness: After a grounded action is executed, the evaluator measures downstream performance under the given task specification (e.g., F1 for classification and 1-RAE for regression). Dataforge maintains a best-so-far state that records the current best pipeline and its evaluation score. If the new action yields improvement beyond a small tolerance, the best state is updated and the action trajectory is stored as a positive experience; otherwise, the system rolls back to the best-so-far pipeline and triggers re-planning to explore alternative actions. This loop enables Dataforge to iteratively search over feature operations without human intervention, while remaining robust to noisy or non-monotonic improvements.
Stopping and budget control: Both loops are governed by user-defined budgets and early-stop criteria (e.g., maximum actions, minimum improvement, and time budget). The optimization terminates when the budget is exhausted or when no meaningful improvement is observed for several iterations, returning the best validated pipeline and its transformed dataset. Overall, the two loops form a closed-cycle mechanism that couples safe executability with evaluation-driven adaptivity.
3.5. Automatic Summary and Reporting
@@
To ensure transparency and provenance, the system automatically logs the routing choices, all actions, grounding outcomes, machine learning performance metrics, and summarizes these logs into a structured, concise natural language report that outlines: (i) the cleaning and feature operations applied, (ii) the evolution of routing decisions, (iii) the validation checks performed, and (iv) the performance progression across iterations. The summary, viewable through the demo interface, translates internal reasoning into a human-readable narrative that lets non-experts trace data transformations and decisions, but also allows the feedback loop to collect experiences as benchmark data and perform self-reflection.
3.6. Pseudo-code of Dataforge
Algorithm 1 Dataforge (Compressed)
0:  Dataset $\mathcal{D}$; budget $\mathcal{B}$ (MaxActions, TimeBudget, MinImprove)
0:  Best transformed data $\mathcal{D}^{\star}$ and pipeline $\pi^{\star}$
1:  $\mathcal{D}^{\star}\leftarrow\textsc{Clean}(\mathcal{D})$; $\pi^{\star}\leftarrow[\,]$; $u^{\star}\leftarrow$-$\infty$
2:  for $t=1$ to MaxActions do
3:   if $\textsc{Stop}(\mathcal{B})$ then
4:    break
5:   end if
6:   $c\leftarrow\textsc{CoarseRoute}(\mathcal{D}^{\star})$ {Sel/Trans/Eng}
7:   $a\leftarrow\textsc{PlanAction}(c,\mathcal{D}^{\star})$;
8:   if $\neg\,\textsc{Ground}(a,\mathcal{D}^{\star})$ then
9:    continue
10:   end if
11:   $\mathcal{D}^{\prime}\leftarrow\textsc{Execute}(a,\mathcal{D}^{\star})$; $u\leftarrow\textsc{Evaluate}(\mathcal{D}^{\prime})$
12:   if $u>u^{\star}+\textsc{MinImprove}(\mathcal{B})$ then
13:    $\mathcal{D}^{\star}\leftarrow\mathcal{D}^{\prime}$; $\pi^{\star}\leftarrow\pi^{\star}\circ[a]$; $u^{\star}\leftarrow u$
14:   end if
15:  end for
16:  return $(\mathcal{D}^{\star},\pi^{\star})$
Algorithm 1 summarizes the overall workflow of Dataforge. Starting from a cleaned dataset, the agent iteratively (i) performs coarse routing to choose a feature-operation family, (ii) plans a concrete action, (iii) grounds the action to ensure executability, and (iv) executes and evaluates the updated data. The best-so-far pipeline is updated only when the downstream performance improves beyond a threshold, and the process terminates by budget or auto-stop.